OpenStudy (anonymous):

what is the sum and product of the roots of each quadratic of x^2-2x-3=0?

OpenStudy (anonymous):

x^2-2x-3=0 compare it with ax^2 + bx+c = 0 we get a=1, b=-2, c=-3 sum of roots = -b/a = -(-2)/1= 2 product of roots= c/a = -3/1 = -3